OpenStudy (goldeverything):

Evaluate each expression for u = –5 and v = –2. Match Term Definition two u minus eight plus three v A) twenty-two seven v minus u plus four B) negative twenty-four eleven minus three v minus u C) negative five nine plus u minus D) 6

OpenStudy (goldeverything):

@e.cociuba can you help me please

OpenStudy (e.cociuba):

Okay so can you rearrange this question. It's a bit confusing. Not organized :)

OpenStudy (anonymous):

"two u mines eight plus three v" means : \[2u-8+3v=2\times(-5)-8+3\times(-2)=-10-8-6=-24\] It is very simple. The others can be done by the same way !

OpenStudy (goldeverything):

i typed it organised but it came out unorganized

OpenStudy (goldeverything):

thank you @Noura11 if you wouldn't mind could you do the other 3 please :)